OpenStudy (anonymous):

What is the function rewritten in vertex form? y = x^2 - 4

OpenStudy (anonymous):

\[(x-0)^{2}=y+4\]

OpenStudy (callisto):

@alexis_holmes Don't close the question and re-post it again. You can, instead, click the 'bump' button.

OpenStudy (anonymous):

(0,-4) vertex

OpenStudy (anonymous):

@Callisto well whats the answer? my bump button wasnt working

OpenStudy (callisto):

@alexis_holmes Please report this problem (bump button is not working) by starting a post in the OpenStudy Feedback group. Also, as I've mentioned in your previous post, *DON'T* just ask for answer.

OpenStudy (anonymous):

k. well your the only one who has a problem with it so flutter off.

OpenStudy (anonymous):

flutter off

OpenStudy (anonymous):

Basically vertex formula is given by: \[\color{green}{y = (x-h)^2 + k}\] Compare this one with: \[y = (x - 0)^2 + (-4)\] Find h and k by comparing the two...

OpenStudy (anonymous):

I just need the answr I tried this question a million times

OpenStudy (callisto):

@alexis_holmes Do respect others. Simply looking for answers violates the code of conduct of this site. Please read it at http://www.openstudy.com/code-of-conduct

OpenStudy (anonymous):

k bye now